# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v1

@@ Abstract
Satellites are increasingly adopting on-board Artificial Intelligence (AI) techniques to enhance platforms’ autonomy through edge inference. In this context, the utilization of deep learning (DL) techniques for segmentation in HS satellite imagery offers advantages for remote sensing applications, and therefore, we train 16 different models, whose codes are made available through our study, which we consider to be relevant for on-board multi-class segmentation of HS imagery, focusing on classifying oceanic (sea), terrestrial (land), and cloud formations. We employ the HYPSO-1 mission as an illustrative case for sea-land-cloud segmentation, and to demonstrate the utility of the segments, we introduce a novel sea-land-cloud ranking application scenario. Our system prioritizes HS image downlink based on sea, land, and cloud coverage levels from the segmented images. We comparatively evaluate the models for in-orbit deployment, considering performance, parameter count, and inference time. The models include both shallow and deep models, and after we propose four new DL models, we demonstrate that segmenting single spectral signatures (1D) outperforms 3D data processing comprising both spectral (1D) and spatial (2D) contexts. We conclude that our lightweight DL model, called 1D-Justo-LiuNet, consistently surpasses state-of-the-art models for sea-land-cloud segmentation, such as U-Net and its variations, in terms of performance (0.93 accuracy) and parameter count (4,563). However, the 1D models present longer inference time (15s) in the tested processing architecture, which is clearly suboptimal. Finally, after demonstrating that in-orbit image segmentation should occur post L1b radiance calibration rather than on raw data, we additionally show that reducing spectral channels down to 3 lowers models’ parameters and inference time, at the cost of weaker segmentation performance.
@@ 1 Introduction
Artificial Intelligence (AI) has seen a growing utilization aboard satellite platforms for remote sensing, benefiting missions equipped with various technologies from conventional RGB to both Multispectral Imaging (MSI) and Hyperspectral Imaging (HSI). The importance of monitoring the Earth’s environment with HSI technology is emphasized by the use of hyperspectral (HS) imagers in numerous contemporary satellite missions, e.g., ESA/Phi-Sat-1 [1], ASI/PRISMA [2, 3], DLR/EnMAP [4], and NTNU/HYPSO-1 [5]. The importance of HS imagers on satellites is further highlighted by upcoming missions, e.g., NASA/PACE, GLIMR and SBG [6], ESA/CHIME [7], ASI/PRISMA Second Gen [8], as well as NTNU/HYPSO-2 and NTNU/HYPSO-3 [9]. HS missions offer great potential due to their detailed spectral characterization of Earth’s features, but they also share a common limitation with other optical remote sensing technologies: the challenge of cloud formation in the atmosphere obstructing the satellites’ capacity to monitor targets of interest. Techniques such as semantic segmentation proves valuable in a broad range of space applications, e.g., from cloud recognition to wildfire detection [10, 11, 12, 13, 14, 15]. Segmentation can indeed serve as a cloud detection technique, categorizing image pixels as either cloudy or non-cloudy, essentially performing binary classification. As a mission example, after the launch in 2020 of the Phi-Sat-1 HS mission part of ESA’s EOP initiative [1], a significant milestone was marked in 2021 when, for the first time, on-board AI was introduced in a satellite for object detection, with the mission’s initial focus being the detection of clouds aboard the satellite. Phi-Sat-1 marked the first in-orbit inference with its convolutional CloudScout Deep Learning (DL) model for cloud detection [10, 13]. The mission’s imager, HyperScout-2, captures the Visible and Near-Infrared (VNIR) and Thermal Infrared (TIR) spectral ranges, and subsequently deploys data processing via AI models in its on-board configurable platform facilitating updates for its algorithms in flight [1, 11]. These in-orbit updates rely on an embedded configurable Commercial off-the-shelf (COTS) platform, which is based on a Vision Processing Unit (VPU) [11, 16]. Furthermore, beyond the Phi-Sat-1 mission, the literature [12] also extends cloud segmentation to images of the missions Sentinel-2B (MSI) and FACSAT-1 (RGB) [17]. For instance, the state of the art [12] deploys a satellite data priority system to compute the cloud coverage level via DL-based segmentation on microcontrollers (MCUs), which determines the image downlink priority. Furthermore, the literature demonstrates that cloud detection can enhance data management aboard satellite platforms as presented next. On the one hand, selective HSI compression methods may be applied to cloud segments via CCSDS-123.0-B-2 to reduce their encoding rate [18, 19], while methods like Compressed Sensing (CS) eliminate on-board compression of cloudy data, shifting computation to ground-based reconstruction [20, 21]. In summary, segmentation can lead to selective compression, reducing the bit rate of data segregated as cloudy or its absolute removal by data screening [11]. Consequently, this streamlines postprocessing and enables the satellite to downlink only valuable non-cloudy data, thus optimizing the efficient usage of the usually limited communication channel [12]. On the other hand, as AI-driven cloud detection segregates non-cloud data from cloudy data, on-board cloud segmentation can additionally play a role in reducing decision-making latency, increasing satellite autonomy with more optimal on-board operations [11, 12, 14, 22, 23, 24, 25]. Furthermore, from a broader perspective, the same models initially trained for cloud binary segmentation can be retrained to solve more complex problems, e.g., multi-class segmentation, where each image pixel is assigned to one of several categories. To further illustrate the importance of AI-powered cloud and multi-class segmentation for many applications, we next employ a particular small satellite mission as an example case. Namely, for space initiatives such as HYPSO-1 (HYPerspectral small Satellite for Ocean observation) [5], the reduction in latency is especially important for time-sensitive remote sensing applications, such as early event warnings. HYPSO-1 focuses on detecting marine phenomena such as Harmful Algal Blooms (HABs), which can lead to significant fish farm salmon mortality. To this extent, segmentation for adequate on-board data management can play a crucial role in achieving prompt management responses to mitigate adverse effects in such scenarios. In the context of multi-class segmentation, by segregating the data along various categories of interest, on-board AI can properly enable the activation of other agents in observational hierarchies monitoring the environment at multiple spatial and temporal scales with drones and marine vessels [5, 26, 27], where the HYPSO-1 mission serves as the space segment of the observational pyramid. However, handling increasingly advanced tasks on satellites, such as segmentation of growing complexity, is feasible only if COTS are conveniently optimized. These hardware solutions provide benefits by accelerating, with high power efficiency, the repetitive operations commonly found in DL networks [1, 13, 16, 22]. Indeed, to enable AI applications in satellites, the literature explores various COTS hardware from the conventional MCUs [12] to the high-performance VPUs aboard Phi-Sat-1, Field-Programmable Gate Arrays (FPGAs) on board HYPSO-1, Tensor Processing Units (TPUs), Graphics Processing Units (GPUs), and other dedicated accelerators [1, 9, 10, 28, 29, 30].
@@
To our knowledge, the existing literature in the field of cloud detection on board satellite platforms does not compare how different shallow and deep learning models, including distinctions like 1D vs. 2D-CNN-based models, influence the performance and complexity of cloud segmentation. Furthermore, the field also overlooks the effects of reducing the number of spectral channels, particularly in MSI/HSI data, on models for cloud segmentation. The existing literature lacks comparisons of key aspects, such as whether on-board AI for cloud detection, is better applied to raw data or L1b radiance data. It does not discuss how calibration impacts segmentation results and does not provide justification for the optimal stage in the on-board data processing pipeline for image segmentation. Moreover, within the literature pertaining to cloud detection in HS satellite imagery, the primary emphasis naturally lies on binary segmentation, specifically distinguishing between the presence of clouds and non-cloudy data; in fact, as earlier mentioned, there are ranking systems that assess cloud levels based on the segmented outcomes. Nevertheless, none of the current systems perform multi-class semantic segmentation for the on-board detection of clouds, sea, and land categories, which typically account for most of the variance in HS satellite imagery. The absence of sea-land-cloud segmentation systems impedes, as a result, the development of applications like priority ranking systems that can integrate these three categories into a unified ranker. In this work, we tackle all the mentioned research gaps and utilize the HYPSO-1 mission as a case study, using “The HYPSO-1 Sea-Land-Cloud-Labeled Dataset” openly accessible in our prior article [31]. Following a methodology inspired by the data science CRISP-DM approach (CRoss Industry Standard Process for Data Mining) to ensure robust and reliable results, we train and compare 16 different Machine Learning (ML) and DL models. This extends the initial model baseline for HYPSO-1 established in [31], and our research findings make contributions to the field of on-board AI in the following manners:
@@
    We propose and fine-tune four new 1D-CNN and 2D-CNN models, with the aim of adapting them for on-board deployment on satellites. This optimization entails reducing the number of model parameters while maintaining performance.
@@
    We demonstrate that 1D-CNN models exhibit superior performance when compared to 2D-CNNs. Notably, our optimized 1D-CNN models like 1D-Justo-LiuNet and 1D-Justo-HuNet consistently outperform other state-of-the-art models, particularly in the case context of multi-class sea-land-cloud HS segmentation for the HYPSO-1 mission.
@@
    We substantiate that segmenting L1b radiance data, as opposed to using raw data, enhances the segmentation of the top-performing models.
@@
    Finally, we propose an image ranking system designed for downlink prioritization based on sea, land, and cloud coverage.
@@ 2.1 Data Source: A Hyperspectral Labeled Datastet from HYPSO-1
We employ “The Sea-Land-Cloud-Labeled Dataset” [31], publicly accessible at https://ntnu-smallsat-lab.github.io/hypso1_sea_land_clouds_dataset. The dataset comprises pixel-level labels for 25 million pixels across 38 scenes captured by HYPSO-1, where each HS image has spatial dimension of 956 x 684 pixels. Each pixel, with ground resolution of approximately 100 m x 600 m in the across- and along-track directions [5, 26], contains a spectral signature of 120 channels covering the 400 to 800 nm range in the VNIR spectrum, with 5 nm spectral resolution. The dataset consists of diverse scenes from various environments and provides raw data and L1b calibrated radiance for comparison with other HS satellites. As for the ground-truth labels, the dataset includes expert pixel-level annotations for three classes: “sea”, “land”, and “clouds”, which are basic categories due to the lack of on-site validation impeding the reliable labeling of more detailed oceanic and terrestrial subcategories. The dataset exhibits a slight class imbalance with 37.01% pixels representing sea, 40.14% representing land, and 22.85% representing clouds/overexposed areas.
2.2 Data Preprocessing: Dimensionality Reduction and Normalization
While our primary focus is HSI segmentation, rather than dimensionality reduction extensively studied for HSI [32], we still conduct spectral channel reduction as a data preprocessing step prior to training ML and DL models. This channel reduction is crucial when dealing with high-dimensional HSI data, as not all channels may contain valuable information. The method in this section helps identify bands for potential removal from the sea-land-cloud dataset. After dimensionality reduction, we apply min-max normalization to maintain a consistent data range and address any potential outlier effects.
Our channel reduction process begins with calculating the statistical standard deviation for each channel (L1b radiance), which provides an initial insight into potential unusual channels such as those with exceptionally low standard deviations demanding further investigation. In particular, we follow two approaches for detecting the most divergent points within the set of deviations shown in Fig. 1: visual inspection for potential outliers (e.g. extremely low deviation values), and the use of algorithms like the Isolation Forest (IF), which consists of an anomaly detection technique to assist identifying anomalous deviation points.
The IF algorithm employs a sensitivity threshold empirically set at 0.08 to balance anomaly detection. In Fig. 1, the red crosses indicate anomalous points detected by the algorithm, specifically flagging the first four standard deviations at $\lambda$=387.85, 391.40, 394.96, and 398.51 nm (blue spectra) as anomalies. The algorithm assists finding out that these four channels actually have zero standard deviations as all their data samples are equal to zero. We observe that this phenomenon is exclusive to L1b radiance and does not occur in the raw data. This suggests that it is a result of the radiometric calibration procedures applied in [31]. Additionally, Fig. 1 reveals that across the dataset, there is a decline in standard deviation values near $\lambda$=760 nm (NIR), primarily due to a sharp reduction in captured light intensity around this wavelength. This decline in light intensity is consistently observed across the image pixels in the dataset. The phenomenon is often associated with atmospheric factors, such as the significant absorption of light by oxygen at this wavelength. The IF algorithm also identifies deviations around $\lambda$=760 nm as anomalous. In particular, it identifies as anomalies the deviations at $\lambda$=734.03, 744.34, 747.77, 758.06, 761.49, and 768.34 nm (channels 99, 102, 103, 106, 107, and 109, starting from 0). Some of these deviations are within the descent peak around $\lambda$=760 nm, while others are just outside, likely due to the 0.08 sensitivity threshold mentioned earlier. In any case, we conclude that channels around the decline at $\lambda$=760 nm may not provide valuable information for pixel-level segmentation. Furthermore, we also add channel 108 ($\lambda$=764.91 nm), as it falls within the descent peak but its deviation remains undetected by the IF algorithm. Consequently, we consider that only the following NIR channels potentially lack meaningful information for segmentation: $\lambda$=758.06, 761.49, 764.91, and 768.34 nm (channels 106, 107, 108, and 109).
Figure 1: Spectral deviations across channels (L1b radiance).
In summary, we suggest removing eight channels from the sea-land-cloud dataset (L1b radiance): the first four with data samples equal to 0 at $\lambda$=387.85, 391.40, 394.96, and 398.51 nm (blue) and the remaining four with a decline at $\lambda$=758.06, 761.49, 764.91, and 768.34 nm (NIR). This reduces the dataset from its original 120 channels to 112 channels, which we utilize in training our ML and DL models in the subsequent methodology. We opt not to exclude further additional channels, as we exercise prudence in our selection to mitigate the risk of inadvertently excluding channels that could have relevance for segmentation. We acknowledge that other methods exist for more accurately identifying channels with valuable information [33, 32], but this falls outside the scope of our current work, which is primarily focused on on-board segmentation.
Since this work additionally studies the impact on the segmentation outcomes when further reducing the number of channels [33], for the particular context of sea-land-cloud HSI segmentation, we utilize the 112-channel data to generate another dataset with significantly fewer channels. Considering that the legacy algorithms for cloud detection in HS satellite imagery typically select 3 channels to mimic RGB image processing [10, 11, 12, 13], we create a new dataset from the original 112 channels, reducing it to merely 3 channels. To achieve this reduction from 112 to 3 channels, we explore the extensive literature on HSI dimensionality reduction techniques, which includes methods like Principal Component Analysis (PCA) [32] and Independent Component Analysis (ICA) [34]. Upon the normalization of the dataset and in alignment with these approaches, we employ PCA to reduce the 112 channels to 3, similarly as in RGB. However, to alleviate to a certain extent the computational demands of satellites operating with tight power constraints, we propose to conduct both the training and inference of PCA exclusively on the ground. Eliminating PCA inference on the satellite enables potential power savings, which could be essential to conduct the subsequent segmentation inference during flight. Consequently, and as also outlined by the literature [11], we suggest implementing on board a direct channel selection (3 bands) instead of deploying PCA inference. In this work, the purpose of employing PCA on ground is not to obtain a low-dimensional dataset directly with the PCA technique, but rather to serve as a tool for identifying the channels containing the most valuable information. In particular, we identify through PCA, for a single principal component, the channels that yield the most substantial influence on the component during the image reduction process. The influence of each channel in the principal component is measured via the component weights (one loading per channel), indicating the particular extent of each channel’s contribution to the component. In the case of the sea-land-cloud dataset (112 channels in L1b radiance), we find that the majority of influential channels are predominantly within the NIR range. However, given the necessity of considering other spectral ranges in addition to NIR for the adequate segmentation of sea, land, and clouds – as the NIR range alone is likely to be insufficient – we opt to pick the top three influential channels (with highest weight loadings) from only the visible blue, second the visible green and red combined, and finally the NIR. In the dataset, the most influential channels within each spectral range are: 412.72 nm (channel 7) for the blue spectrum, 699.61 nm (channel 89) in the green and red spectra, and 747.77 nm (channel 103) in the NIR spectrum. These align closely with the L1b radiance bands frequently analyzed in HYPSO-1, especially for channels 89 and 103.
2.3 Semantic Segmentation: Machine Learning and Deep Learning Models

### v2

@@ Abstract
Satellites are increasingly adopting on-board AI for enhanced autonomy through in-orbit inference. In this context, the use of deep learning (DL) techniques for segmentation in hyperspectral (HS) satellite imagery offers advantages for remote sensing applications, and therefore, we train 16 different models, whose codes are made available through our study, which we consider to be relevant for on-board multi-class segmentation of HS imagery, focusing on classifying oceanic (sea), terrestrial (land), and cloud formations. We employ the HYPSO-1 mission as an illustrative case for sea-land-cloud segmentation, and to demonstrate the utility of the segments, we introduce a novel sea-land-cloud ranking application scenario. We consider how to prioritize HS image downlink based on sea, land, and cloud coverage levels from the segmented images. We comparatively evaluate the models for future in-orbit deployment, considering performance, parameter count, and inference time. The models include both shallow and deep models, and after we propose four new DL models, we demonstrate that segmenting single spectral signatures (1D) outperforms 3D data processing comprising both spectral (1D) and spatial (2D) contexts. We conclude that our lightweight DL model, called 1D-Justo-LiuNet, consistently surpasses state-of-the-art models for sea-land-cloud segmentation, such as U-Net and its variations, in terms of performance (0.93 accuracy) and parameter count (4,563). However, the 1D models present longer inference time (15s) in the tested processing architecture, which seems to be a suboptimal architecture for this purpose. Finally, after demonstrating that in-orbit segmentation should occur post L1b radiance calibration rather than on raw data, we also show that reducing spectral channels down to 3 lowers models’ parameter counts and inference time, at the cost of weaker segmentation performance.
@@ 1 Introduction
Artificial Intelligence (AI) has seen a growing utilization aboard satellite platforms for remote sensing, benefiting Earth observation missions equipped with various technologies from conventional RGB to both Multispectral Imaging (MSI) and Hyperspectral Imaging (HSI). The importance of monitoring the Earth’s environment with HSI technology is emphasized by the use of hyperspectral (HS) imagers in numerous contemporary satellite missions, e.g., ESA/Phi-Sat-1 [1], ASI/PRISMA [2, 3], DLR/EnMAP [4], and NTNU/HYPSO-1 [5]. The importance of HS imagers on satellites is further highlighted by upcoming missions, e.g., ESA/CHIME [6, 7], NASA/PACE, GLIMR and SBG [8], ASI/PRISMA Second Gen [9], as well as NTNU/HYPSO-2 and NTNU/HYPSO-3 [10]. HS missions offer great potential due to their detailed spectral characterization of Earth’s features, but they also share a common limitation with other optical remote sensing technologies: the challenge of cloud formation in the atmosphere obstructing the satellites’ capacity to monitor targets of interest. In this context, techniques such as semantic segmentation proves valuable in a broad range of space applications, e.g., from cloud recognition to wildfire detection [11, 12, 13, 14, 15, 16, 17, 18]. Segmentation can indeed serve as a cloud detection technique, categorizing image pixels as either cloudy or non-cloudy, essentially performing binary classification. As a mission example, after the launch in 2020 of the Phi-Sat-1 HS mission part of ESA’s EOP initiative [1], a significant milestone was marked in 2021 when, for the first time, on-board AI was introduced in a satellite for object detection, with the initial focus being the detection of clouds aboard the satellite. Phi-Sat-1 marked the first in-orbit inference with its convolutional CloudScout Deep Learning (DL) model for cloud detection [11, 14]. The mission’s imager, HyperScout-2, captures the Visible and Near-Infrared (VNIR) and Thermal Infrared (TIR) spectral ranges, and subsequently deploys data processing via AI models in its on-board configurable platform facilitating updates for its algorithms in flight [1, 12]. These in-orbit updates rely on an embedded configurable Commercial off-the-shelf (COTS) platform, which is based on a Vision Processing Unit (VPU) [12, 19]. Furthermore, beyond the Phi-Sat-1 mission, the literature [13] also extends cloud segmentation to image data of the missions Sentinel-2B (MSI) and FACSAT-1 (RGB) [20]. For instance, the state of the art [13] deploys a satellite data priority system to compute the cloud coverage level via DL-based segmentation on microcontrollers (MCUs), which determines the image downlink priority. Furthermore, the literature demonstrates that cloud detection can enhance data management aboard satellite platforms as presented next. On the one hand, selective HSI compression methods may be applied to cloud segments via CCSDS-123.0-B-2 to reduce their encoding rate [21, 22], while other methods like Compressed Sensing (CS) can also be employed to handle cloudy data, where CS proves particularly advantageous for satellites as it transfers computational complexity from orbit to mere ground-based reconstruction [23, 24]. Therefore, segmentation can lead to selective compression reducing the bit rate of data segregated as cloudy, or its absolute removal by data screening [12]. Consequently, this streamlines data postprocessing and enables the satellite to downlink, for instance, only valuable non-cloudy data, thus optimizing the usage of the often limited communication channel [13]. On the other hand, on-board cloud segmentation can additionally play a role in reducing decision-making latency, increasing satellite autonomy with more optimal on-board operations [12, 13, 15, 25, 26, 27, 28]. Furthermore, from a broader perspective, the models first trained for cloud binary segmentation can be trained again to solve more complex problems instead, e.g., multi-class segmentation, where each image pixel is assigned to one of various categories. To further illustrate the importance of AI-powered cloud and multi-class segmentation for space applications, we next employ a particular small satellite mission as an example case. Namely, for Earth observation initiatives such as HYPSO-1 (HYPerspectral small Satellite for Ocean observation) [5], the reduction in decision-making latency is especially important for time-sensitive remote sensing applications, such as early event warnings. HYPSO-1 focuses on timely detection of marine phenomena such as Harmful Algal Blooms (HABs), which can lead to significant mortality of farmed salmon. To this extent, in-orbit segmentation for adequate on-board data management can play a crucial role in achieving prompt management responses to mitigate adverse effects in such scenarios. In the context of multi-class segmentation, by segregating the data along various categories of interest, on-board AI can activate other subsequent agents in observational hierarchies monitoring the environment at multiple spatial, spectral and temporal scales with e.g. drones and marine vessels [5, 29, 30], where the HYPSO-1 satellite may serve as the space segment of an observational pyramid. However, handling increasingly advanced tasks on satellites, such as segmentation of growing complexity, is feasible only when the COTS components within the satellite’s payload are appropriately designed and optimized. Some COTS platforms provide benefits by accelerating, with high power efficiency, the repetitive operations commonly found in DL networks [1, 14, 19, 25]. Indeed, to enable AI applications in satellites, the literature explores various COTS hardware from the conventional MCUs [13] to the high-performance VPUs aboard Phi-Sat-1, Field-Programmable Gate Arrays (FPGAs) on board HYPSO-1, Tensor Processing Units (TPUs), Graphics Processing Units (GPUs), and other dedicated accelerators [1, 10, 11, 31, 32, 33].
@@
To our knowledge, the existing literature in the field of cloud detection on board satellite platforms does not compare how different shallow and deep learning models, including distinctions like 1D vs. 2D-CNN-based models, influence the performance and complexity of cloud segmentation. Furthermore, the field also overlooks the effects of reducing the number of spectral channels, particularly in MSI/HSI data, on models for cloud segmentation. The existing literature lacks comparisons of key aspects, such as whether on-board AI for cloud detection, is better applied to raw data or L1b calibrated radiance data. It does not discuss how L1b calibration impacts segmentation results and does not provide justification for the optimal stage in the on-board data processing pipeline for image segmentation. Moreover, within the literature pertaining to cloud detection in HS satellite imagery, the primary emphasis naturally lies on binary segmentation, specifically distinguishing between the presence of clouds and non-cloudy data; in fact, as earlier mentioned, there are ranking systems that assess cloud levels based on the segmented outcomes. Nevertheless, none of the systems at the time of writing perform multi-class semantic segmentation for the simultaneous detection of clouds, sea, and land categories, which typically account for most of the variance in HS satellite imagery. The absence of sea-land-cloud segmentation systems impedes, as a result, the development of applications like priority ranking systems that can integrate the three categories into a unified ranker. In this work, we tackle the mentioned research gaps and utilize the HYPSO-1 mission as a case study, using its “Sea-Land-Cloud-Labeled Dataset” openly accessible in our prior article [34] (see data availability section at the end of this paper).
Following a methodology inspired by the data science CRISP-DM approach (CRoss Industry Standard Process for Data Mining) to ensure robust and reliable results, we train and compare 16 different Machine Learning (ML) and DL models. This extends the segmentation model baseline for HYPSO-1 first established in [34], and our research findings make contributions to the field as follows:
@@
    We propose and fine-tune four new 1D-CNN and 2D-CNN models, with the aim of adapting them for on-board deployment on satellites. This optimization entails reducing the number of model parameters to achieve highly compact lightweight models while maintaining performance.
@@
    For the case context of multi-class sea-land-cloud HS segmentation for the HYPSO-1 mission, we demonstrate that 1D-CNN models exhibit superior performance when compared to 2D-CNNs. Indeed, our optimized 1D-CNN lightweight models, 1D-Justo-LiuNet and 1D-Justo-HuNet, consistently outperform other state-of-the-art models.
@@
    We substantiate that segmenting L1b radiance data, as opposed to using raw sensor data, enhances the segmentation of the top-performing models.
@@
    Finally, we propose a case example consisting of an image ranking system intended for downlink prioritization based on sea, land, and cloud coverage.
@@ 2.1 Data Source: A Hyperspectral Labeled Datastet from HYPSO-1
We employ the “Sea-Land-Cloud-Labeled Dataset” from HYPSO-1 [34] (see data availability section). The dataset consists of pixel-level labels for 25 million pixels across 38 scenes captured by HYPSO-1, where each HS image has spatial dimension of 956 $\times$ 684 pixels. Each pixel, contains a spectral signature of 120 channels covering the 400 to 800 nm range in the VNIR spectrum, with approximately 5 nm spectral resolution [5, 29]. The dataset consists of diverse scenes from various environments and provides raw data and L1b calibrated radiance for comparison with other HS missions. The dataset includes expert pixel-level annotations for three classes: “sea”, “land”, and “clouds”. The dataset exhibits a slight class imbalance with 37.01% pixels representing sea, 40.14% representing land, and 22.85% representing clouds/overexposed areas.
2.2 Channel Reduction Preprocessing
Following the common practice to inspect the data features prior to training ML and DL models, we next identify any unusual channel features that may require correction. In the sea-land-cloud dataset [34], we observe that the first four channels with wavelengths $\lambda$=388, 391, 395, and 399 nm (visible blue spectrum) exhibit data samples equal to 0 for all the dataset HS images. We observe that this phenomenon occurs only to the L1b radiance, however, it is not present in the raw sensor data. This suggests that the occurrence of zero values in the first channels is a result of the radiometric calibration procedures applied by [34] following [35]. Furthermore, we observe that the dataset, near its wavelength $\lambda$=760 nm (oxygen A-band within the NIR spectrum), presents a sharp decline in light intensity which is attributed to the significant light absorption by the atmosphere’s oxygen at this wavelength [36, 37]. In summary, we exclude first wavelengths $\lambda$=388, 391, 395, and 399 nm, and we also exclude the wavelengths in close proximity to $\lambda$=760 nm, namely, $\lambda$=758, 761, 765, and 768 nm. As a result, the dataset’s original 120 channels are reduced to 112. The 112 channels, after undergoing min-max normalization, are the ones we employ for training the ML and DL models in the subsequent methodology. Since this work additionally studies the impact on the segmentation outcomes when significantly reducing the number of channels for the particular context of sea-land-cloud HSI segmentation, we use the 112-channel data to create a new dataset containing only 3 channels as detailed in A.
2.3 Segmentation Modeling